# Deep Submodular Peripteral Networks

### v2

@@ Appendix F Analysis of Peripteral Loss
We first plot the loss and the gradients in fig. 7. From the plots in fig. 7 and gradient in eq. 23, it is clear that if ${\Delta}$ is (very) positive, then ${\delta}$ must also be (very) positive to produce a small loss. If ${\Delta}$ is (very) negative then so also must ${\delta}$ be to produce a small loss. The absolute value of ${\Delta}$ determines how non-smooth loss is so we include an outer pre-multiplication by $|{\Delta}|$ ensures gradient of loss, as a function of ${\delta}$, asymptotes to negative one (or one when ${\Delta}$ is negative) in the penalty region, analogous to hinge loss, and thus produces numerically stable gradients.
Having established the “mother” loss $\mathcal{L}_{\Delta}({\delta})$ we can see that ${\Delta}$ controls its curvature (the smaller the $|{\Delta}|$ the higher the curvature). However, for each example, this margin would be different, and therefore, to provide global control over the curvature of the loss, we smooth it out using $\beta$ in a way that produces smoother, more stable, and convergent optimization. This leads to the following loss (in addition to the explanation provided in the section 3)-
@@
and the corresponding gradient will be -
$\frac{\partial L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}=-\operatorname{sgn}({\Delta})\sigma\left(\beta\left(\tau-\frac{{\delta}}{\kappa{\Delta}+\varepsilon\operatorname{sgn}{{\Delta}}}\right)\right)$ (25)
We plot the gradients in fig. 9 for two offsets, namely $\tau=1$ and $\tau=5$. From these plots, it is clear that $\beta$ provides control over the smoothness of $L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})$ where the smaller the $\beta$ the smaller the curvature (without impacting the gradient magnitudes). The importance of low curvature of the loss objective has been studied from the optimization stability [Gilmer et al., 2021] as well as from the lens of robustness in deep learning [Srinivas et al., 2022, Foret et al., 2021]. Similarly, we can also see that $\tau$ merely provides an offset in the gradient, which, however, can nevertheless be important for the gradient magnitude while we are optimizing.
Lastly, to account for corner cases where ${\Delta}$ is very small,
$L_{{\Delta};\alpha,\beta,\tau,\kappa,\varepsilon}({\delta})=\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))|{\delta}|+(1-\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\mathcal{L}_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})$ (26)
and the corresponding gradient will be -
$\frac{\partial L_{{\Delta};\alpha,\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}=\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\operatorname{sgn}({\delta})+(1-\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\frac{\partial L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}$ (27)
Note that for ${\delta}=0$ the function is indeed non-differentiable, however, we can have any subgradient between $[-1,1]$. Similar to PyTorch [Paszke et al., 2019] we choose to have the subgradient to be 0, as $|{\delta}|$ is also 0 at that point. In practice, we do not encounter any instabilities due to this, and gate value throughout our experiments remains very small.
@@ Appendix G Peripteral Networks and Training Flow
Figure 10: A “peripteral neural temple” showing a Greek/Roman style peripteral structure with many pillars rendered as a form of neural temple, where each of the pillars, the submodular preserving and permutation-invariant aggregation, and the roof are all seen as having an underlying neural network graphical structure.
Figure 11: The figure shows the structure of a DSPN as well as the control flow of how a DSPN is trained.
Figure 11 depicts the DSPN architecture. In fact, we see two copies of a DSPN presented in Figure 11, the left one instantiated for the $E$ set and the right one instantiated for the $M$ set. Each DSPN consists of a set of pillars which transform a number of objects (e.g., images corresponding to a set) into a non-negative embedding space. The reason for the space to be non-negative is to ensure that the subsequent transformations can preserve submodularity, in the same way that a concave composed with a modular function requires the modular function to be non-negative to preserve submodularity. The figure above shows that object space has dimension nine (the number of red nodes on the bottom of each pillar) and embedding space having dimension seven (the seven orange nodes at the top of each pillar). This DSPN has five pillars which means that it is appropriate for an $E$ or $M$ set of size five. The outputs of the pillars are aggregated using a submodular-preserving permutation invariant aggregator (such as a weighted matroid rank function) but in the experiments in this paper, we use a simple uniform matroid which corresponds to a summation aggregator. Note that each scalar output of each pillar is aggregated separately and after aggregation we get a resulting seven-dimensional vector which is input to a deep submodular function (DSF). We consider the aggregation step and the final DSF to be the “roof model.”, highlighted with the blue triangle above. The word peripteral is an adjective (applicable typically to a building) which means having a single row of pillars on all sides in the style of the temples of ancient Greece. This is reminiscent of the above DSPN structures which is the reason for this name.
@@ Appendix H Additional Experimental Details
To assess the efficacy of our peripteral loss, we consider two popular baselines to learn from pairwise comparisons. Note that we only replace our peripteral loss function with the baseline and we do not remove any necessary regularizers such as $\mathcal{L}_{redn}$ and $\mathcal{L}_{aug}$. Using the same notations as the main paper, we define the regression loss $\mathcal{L}_{\text{regr}}$ and $\mathcal{L}_{\text{marg}}$ in equations 28 and 29.
@@
Hyperparameters
We train every DSPN on 2 NVIDIA-A100 GPUs with a cyclic learning rate scheduler  [Smith, 2017], using Adam optimizer  [Kingma and Ba, 2014]. To ensure non-negative parameters in DSF, we perform projected gradient descent, where we apply projection on ${\psi_{{w_{\mathfrak{r}}}}}$ to the non-negative orthant after every gradient descent step.
We list the hyperparameters we use for the peripteral loss in table 2 and table 3. We observed that $\beta$ and $\tau$ are the most salient hyperparameters since they control the curvature and hinge of the loss respectively. The benefits of smaller loss curvature have been widely studied from optimization stability in deep learning  [Gilmer et al., 2021, Srinivas et al., 2022, Foret et al., 2021] and our low value of $\beta$ corroborates with the previous works. $\alpha$ parameter controls the $\tanh$ gate, the smaller the value of alpha the more weightage of our objective function. Since we did not encounter corner cases where ${\Delta}$ is very small, neither $\epsilon$ nor $\alpha$ were found to be salient.
@@
For the results presented in Sections 5.2, we create pathologically imbalanced datasets from sets of images that the DSPN never encountered during training. Specifically, we add duplicates for each image in a class; the number of duplicates varies per class and is drawn from a Zipf distribution  [Zipf, 2013]. We show the final distributions of the imbalanced dataset in Figure 12.
@@ Appendix I Qualitative Analysis of Learnt Features
In Table 1, we show that imparting DSPNs the ability to learn features is crucial for successful training. Feature-based functions and DSFs work best when features correspond to some bag-of-words or count like attribute. For example, Lin and Bilmes [2012], Liu et al. [2013] both use TF-IDF based feature extractors to instantiate feature-based functions. More recently, Lavania and Bilmes [2019] learns “additively contributive" (AC) features with an autoencoder by restricting all weights in the initial layer of the decoder to the positive orthant. Unfortunately, universal features for images procured from modern pretrained models such as CLIP do not necessarily have this property.
We show that DSPNs learn bag-of-words features in Figure 13, by exploring the preferences of individual DSPN/CLIP features. For visualization purposes, we randomly select 10 features from the model in consideration and use each feature to rank the images from a held out set. The top 10 images selected by any of the DSPN features appear very homogeneous and are strongly correlated to class, while images with the lowest feature values (typically with feature value 0) are practically indistinguishable from a random subset of images. In other words, a high value indicates the presence of some attribute while a low value indicates its absence. When we perform a similar qualitative analysis upon CLIP features, we find that both the high valued sets and low valued sets have some weak correlation to class. CLIP features cannot be interpreted in the same way as DSPN because each feature corresponds to a coordinate in some higher dimensional space. Our results show that features that encapsulate count-like attributes are essential for learning a good DSF.

### v3

@@ Appendix F Analysis and Further Description of Peripteral Loss
We first plot the loss and the gradients in Fig. 7. From the plots in Fig. 7 and gradient in Fig. 21, it is clear that if ${\Delta}$ is (very) positive, then ${\delta}$ must also be (very) positive to produce a small loss. If ${\Delta}$ is (very) negative then so also must ${\delta}$ be to produce a small loss. The absolute value of ${\Delta}$ determines how non-smooth loss is so we include an outer pre-multiplication by $|{\Delta}|$ ensures gradient of loss, as a function of ${\delta}$, asymptotes to negative one (or one when ${\Delta}$ is negative) in the penalty region, analogous to hinge loss, and thus produces numerically stable gradients.
Having established the “mother” loss $\mathcal{L}_{\Delta}({\delta})$ we can see that ${\Delta}$ controls its curvature (the smaller the $|{\Delta}|$ the higher the curvature). Still, we augment the mother loss with additional hyperparameters in way that produces smoother and more stable and convergent optimization. Firstly, rather than a requiring a margin of one, we introduce a hyperparameter $\tau$ to express this margin whose default value is $\tau=1$. Second, since the units of the oracle target might be different than the units of the learner, we introduce a unit adjustment parameter $\kappa$ that scales ${\Delta}$. We also introduce an anti-smoothness parameter $\beta>0$ (smaller $\beta$ means smoother loss) to further optionally smooth the transition between the linearly increasing loss region and the zero-loss region of ${\delta}$. That is, since in general the margin could be different from sample to sample, the $\beta$ hyperparameter provides global control over the curvature of the loss in a way that produces smoother, more stable, and convergent optimization. Lastly, if ${\Delta}$ gets small, to avoid finite-precision numerical problems, we add a small signed value $\epsilon\operatorname{sgn}({\Delta})$ for $\varepsilon\geq 0$ which slightly pushes the denominator away from zero depending on the sign of ${\Delta}$. This yields $\mathcal{L}_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})$ which is defined as:
The rate of the gating is determined by $\alpha\geq 0$, and since $\frac{\alpha}{|\kappa{\Delta}|}\geq 0$, the gate switches to the absolute value $|{\delta}|$ whenever ${\Delta}$ gets very small. Thus if the oracle target is indifferent to $E$ vs. $M$ this encourages the learner also to be indifferent. Despite these hyperparameters, the loss gradient is still independent of the margin and $\beta$ making it useful for fine-tuning a learning rate
Considering only following term (which does not have the gating term),
@@
the corresponding gradient will be:
$\frac{\partial L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}=-\operatorname{sgn}({\Delta})\sigma\left(\beta\left(\tau-\frac{{\delta}}{\kappa{\Delta}+\varepsilon\operatorname{sgn}{{\Delta}}}\right)\right).$ (25)
We plot the gradients in fig. 9 for two offsets, namely $\tau=1$ and $\tau=5$. From these plots, it is clear that $\beta$ provides control over the smoothness of $L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})$ where the smaller the $\beta$ the smaller the curvature (without impacting the gradient magnitudes). The importance of low curvature of the loss objective has been studied from the optimization stability [30] as well as from the lens of robustness in deep learning [95, 27]. Similarly, we can also see that $\tau$ merely provides an offset in the gradient, which, however, can nevertheless be important for the gradient magnitude while we are optimizing.
Bringing back in the gating term (again, to account for corner cases where ${\Delta}$ is very small), we have
$L_{{\Delta};\alpha,\beta,\tau,\kappa,\varepsilon}({\delta})=\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))|{\delta}|+(1-\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\mathcal{L}_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta}),$ (26)
and the corresponding gradient will be:
$\frac{\partial L_{{\Delta};\alpha,\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}=\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\operatorname{sgn}({\delta})+(1-\tanh\left(\frac{\alpha}{|\kappa{\Delta}|}\right))\frac{\partial L_{{\Delta};\beta,\tau,\kappa,\varepsilon}({\delta})}{\partial{\delta}}.$ (27)
Note that for ${\delta}=0$ the function is indeed non-differentiable, however, we can have any subgradient between $[-1,1]$. Similar to PyTorch [78] we choose to have the subgradient to be 0, as $|{\delta}|$ is also 0 at that point. In practice, we do not encounter any instabilities due to this, and gate value throughout our experiments remains very small.
@@
Final Extended Loss
Building and extending Section 3.2 using the peripteral loss along with its hyperparameters, and with all of the individual loss components described from Section 3.1, the final total loss $\mathcal{L}_{\text{tot}}(f_{w};{{\Delta}({E}|{M})},{{\delta}_{{f_{w}}}({E}|{M})})$ for an $E,M$ pair becomes:
@@ Appendix G Peripteral Networks and Training Flow
Figure 10: A “peripteral neural temple” showing a Greek/Roman style peripteral structure with many pillars rendered as a form of neural temple, where each of the pillars, the submodular preserving and permutation-invariant aggregation, and the roof are all seen as having an underlying neural network graphical structure. This figure was generated using GPT-4 after an enormous amount of trial and error including many seed images, and then it was further edited in Adobe Photoshop.
Figure 1 in the main paper depicts the DSPN architecture. In fact, we see two copies of a DSPN presented in Figure 1, the left one instantiated for the $E$ set and the right one instantiated for the $M$ set. Each DSPN consists of a set of pillars which transform a number of objects (e.g., images corresponding to a set) into a non-negative embedding space. The reason for the space to be non-negative is to ensure that the subsequent transformations can preserve submodularity, in the same way that a concave composed with a modular function requires the modular function to be non-negative to preserve submodularity. The figure above shows that object space has dimension nine (the number of red nodes on the bottom of each pillar) and embedding space having dimension seven (the seven orange nodes at the top of each pillar). This DSPN has five pillars which means that it is appropriate for an $E$ or $M$ set of size five. The outputs of the pillars are aggregated using a submodular-preserving permutation invariant aggregator (such as a weighted matroid rank function) but in the experiments in this paper, we use a simple uniform matroid which corresponds to a summation aggregator. Note that each scalar output of each pillar is aggregated separately and after aggregation we get a resulting seven-dimensional vector which is input to a deep submodular function (DSF). We consider the aggregation step and the final DSF to be the “roof model.”, highlighted with the blue triangle above. The word peripteral is an adjective (applicable typically to a building) which means having a single row of pillars on all sides in the style of the temples of ancient Greece. This is reminiscent of the above DSPN structures which is the reason for this name (see figure10).
@@ Appendix H Additional Experimental Details
To assess the efficacy of our peripteral loss, we consider two popular baselines to learn from pairwise comparisons. Note that we only replace our peripteral loss function with the baseline and we do not remove any necessary regularizers such as $\mathcal{L}_{redn}$ and $\mathcal{L}_{aug}$. Using the same notations as the main paper, we define the regression loss $\mathcal{L}_{\text{regr}}$ and $\mathcal{L}_{\text{marg}}$ in equations 30 and 31.
@@
Hyperparameters and Sensitivity Analysis
We train every DSPN on 2 NVIDIA-A100 GPUs with a cyclic learning rate scheduler  [94], using Adam optimizer  [47]. To ensure non-negative parameters in DSF, we perform projected gradient descent, where we apply projection on ${\psi_{{w_{\mathfrak{r}}}}}$ to the non-negative orthant after every gradient descent step.
We list the hyperparameters we use for the peripteral loss in table 2 and table 3. We observed that $\beta$ and $\tau$ are the most salient hyperparameters since they control the curvature and hinge of the loss respectively. The benefits of smaller loss curvature have been widely studied from optimization stability in deep learning  [30, 95, 27] and our low value of $\beta$ supports previous work. $\alpha$ parameter controls the $\tanh$ gate, the smaller the value of alpha the more weight is given to the main part of our objective function. Since we did not encounter corner cases where ${\Delta}$ is very small, neither $\epsilon$ nor $\alpha$ were found to be particularly influential.
@@
[…]
k=10 k=20 k=30 k=40 k=50 k=60 k=70 k=80 k=90 k=100
$\beta=0.01$ 0.88 0.89 0.92 0.91 0.91 0.91 0.92 0.92 0.92 0.92
$\beta=0.1$ 0.88 0.88 0.90 0.90 0.90 0.92 0.91 0.92 0.92 0.92
$\beta=0.5$ 0.90 0.89 0.90 0.92 0.92 0.92 0.92 0.93 0.92 0.92
Table 4: Sensitivity of $\beta$ based on normalized FL evaluation.
k=10 k=20 k=30 k=40 k=50 k=60 k=70 k=80 k=90 k=100
$\tau=1$ 0.84 0.84 0.85 0.87 0.88 0.88 0.88 0.89 0.90 0.90
$\tau=5$ 0.90 0.88 0.88 0.89 0.90 0.90 0.90 0.91 0.91 0.92
$\tau=10$ 0.90 0.89 0.90 0.92 0.92 0.92 0.92 0.93 0.92 0.92
Table 5: Sensitivity of $\tau$ based on normalized FL evaluation.
k DSPN DSPN (No Redun) Random
10 0.91 0.88 0.65 $\pm$ 0.04
20 0.93 0.89 0.67 $\pm$ 0.03
30 0.91 0.89 0.67 $\pm$ 0.02
40 0.90 0.88 0.68 $\pm$ 0.02
50 0.90 0.88 0.68 $\pm$ 0.02
60 0.89 0.88 0.68 $\pm$ 0.02
70 0.88 0.88 0.68 $\pm$ 0.02
80 0.88 0.87 0.69 $\pm$ 0.02
90 0.88 0.87 0.69 $\pm$ 0.02
100 0.88 0.87 0.70 $\pm$ 0.02
Table 6: Effect of setting $\lambda_{3}=\lambda_{4}=0$ on normalized FL evaluations on CIFAR100
@@
For the results presented in Sections 5.2, we create pathologically imbalanced datasets from sets of images that the DSPN never encountered during training. Specifically, we add duplicates for each image in a class; the number of duplicates varies per class and is drawn from a Zipf distribution  [116]. We show the final distributions of the imbalanced dataset in Figure 11.
@@ Appendix I Qualitative Analysis of Learnt Features
In Table 1, we show that imparting DSPNs the ability to learn features is crucial for successful training. Feature-based functions and DSFs work best when features correspond to some bag-of-words or count like attribute. For example, [60, 62] both use TF-IDF based feature extractors to instantiate feature-based functions. More recently, [51] learns “additively contributive" (AC) features with an autoencoder by restricting all weights in the initial layer of the decoder to the positive orthant. Unfortunately, universal features for images procured from modern pretrained models such as CLIP do not necessarily have this property.
We show that DSPNs learn bag-of-words features in Figure 12, by exploring the preferences of individual DSPN/CLIP features. For visualization purposes, we randomly select 10 features from the model in consideration and use each feature to rank the images from a completely held out set. The top 10 images selected by any of the DSPN features appear very homogeneous and are strongly correlated to class, while images with the lowest feature values (typically with feature value 0) are practically indistinguishable from a random subset of images. In other words, a high value indicates the presence of some attribute while a low value indicates its absence. When we perform a similar qualitative analysis upon CLIP features, we find that both the high valued sets and low valued sets have some weak correlation to class. CLIP features cannot be interpreted in the same way as DSPN because each feature corresponds to a coordinate in some higher dimensional space. Our results show that features that encapsulate count-like attributes are essential for learning a good DSF.
@@
Appendix J Limitations
A limitation of DSPNs is that they may not be able to produce effective summaries that are significantly larger than the largest sets that they were trained on. In our experiments, we do observe that the learnt DSPN produces high quality summaries of size 500 though they were not trained on sets larger than 100. However, further investigation is needed to determine if DSPNs can continue to produce high quality summaries with significantly larger cardinalities.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims from the abstract and introduction are justified throughout the paper.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Limitations of our paper is discussed in Appendix J.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: We do not make any assumptions in our theoretical results.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We provide all the details required for reproducibility in Appendix H and Table 2.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We plan to release the codebase in a separate GitHub repository under the name bhattg/DeepSubmodularPeripteralNetworks.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: All the details are provided in Appendix H and Table 2.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We provide standard error for the results of all experiments that are non-deterministic in section 5.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We provide the details in Section 5 and Table 2.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We reviewed NeurIPS Code of Ethics and abide by it in our work.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: Our work proposes a new approach to learn submodular functions from an oracle. We do not anticipate any negative social impact of our work.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We do not release any dataset or model.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All datasets we use are cited.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [N/A]
  52. Justification: We do not release any new assets in this work.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: We do not have any experiments with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No human subjects were used on our work.